-> XCIO DE D Question 27 A 150.0-g sample of metal at 80.0°C is added to 150.0 g of H20 at 20.0°C. The temperature rises to 2 solve both pls
Answer #1

Q27:

Heat lost by metal is gained by the water.

m1s1(∆T)1 = m2s2(∆T)2

150.0g×4.18J/g.℃ × (23.3- 20.0)℃ = 150.0g×s2×(80.0-23.3)℃

s2 = 4.18×3.3/56.7 = 0.2433J/g.℃

Answer: 0.24J/(g.℃)

Q28:

Sn + SnO2 + 92kJ \rightarrow 2SnO

That means +92kJ is the heat required to produce 2 mole of SnO. The reaction is endothermic in nature i.e. ∆H > 0.

(1/2)Sn + (1/2)SnO2\rightarrow SnO. ∆H = 92kJ/2 = + 46.kJ

The enthalpy change for production of 1.00mol of SnO = 46kJ

Answer: 46kJ